Sarah is now 3 years older than kevin. 5 years from now she will be 1.25 times as old as Kevin. how old are they now

Answer by edjones(8007)

s=sarah's age, k=kevin's age
s=k+3
s+5=1.25(k+5)
k+3+5=1.25(k+5)Substitute k+3 for s
k+8=1.25k+6.25
-.25k=-1.75
k=7
s=10
.
Ed